Hi can you help with this question:
2m+3=23-8m
isbn is 0618372628

Answer by Nate(3500)

2m + 3 = 23 - 8m
2m + 8m + 3 - 3 = 23 - 3 - 8m + 8m
10m = 20
10m/10 = 20/10
m = 2